OpenStudy (anonymous):

solve for x square root x-4=-1

OpenStudy (anonymous):

there no answer choice for that..

OpenStudy (anonymous):

\[\sqrt{x}-4=-1\] \[\sqrt{x}=-1+4\] \[\sqrt{x}=3\] \[squaring\ on\ both\ sides\] \[=> x=9\]

OpenStudy (anonymous):

the choices are : 25/2 , 1, 5 and no solution

OpenStudy (anonymous):

\[\sqrt{x-4} = -1\] \[squaring\ on\ both\ sides\] \[=> x-4=1\] \[x=1+4=5\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

verifying... @LadyDine , what's your answer?